Karen asks her teacher, Miss Sharp, how old she is.
Miss Sharp replies that if you double her age, add 7 and then divide by 3, you get 21.

Answer by sdmmadam@yahoo.com(530)

Miss Sharp replies that if you double her age, add 7 and then divide by 3, you get 21.
Let Miss Sharp's age b A yrs
Double her age,add 7 and then divide by 3,then you get 21
That is (2A + 7)/3 = 21
Multiplying by 3
(2A+7) = 21X3
2A +7 =63
2A =63-7
2A = 56
A = 56/2 = 28
Answer: Miss Sharp's age is 28 yrs